The length of a rectangle is 5 in. more than the width. If the perimeter is 70in., what is the width?
Answer by stanbon(75887)

The length of a rectangle is 5 in. more than the width. If the perimeter is 70in., what is the width?
------------------
Perimeter = 2(L + W)
70 = 2(W+5 + W)
35 = 2W+5
2W = 30
W = 15" (width)
---
L = W+5 = 20" (length)
============================
Cheers,
Stan H.